\begin{thebibliography}{99}

%\bibitem{Andrews-Curtis}
%{\bf UNUSED REFERENCE:}  \ \ \
%J.~J.~Andrews and M.~L.~Curtis, {\it $n$-Space modulo and arc}, Ann.
%of Math. (2) {\bf 75} (1962) 1--7.

\bibitem{Armentrout}
S. Armentrout,
{\sl A decomposition of $E^3$ into straight arcs and singletons},
Dissertationes Math. Rozprawy Mat. {\bf 68} (1970), 46 pp.

%\bibitem{BDVW}
%M.~Bestvina, R.~J.~Daverman, G.~A.~Venema and J.~J.~Walsh,
%{\it A 4-dimensional 1-LCC shrinking theorem},
%Topol. Appl. {\bf 110} (2001), 3--20.

%\bibitem{bing}
%R.~H.~Bing, {\it The Kline sphere characterization problem}, Bull.
%Amer. Math. Soc. {\bf 52} (1946), 644--653.

%\bibitem{bing1}
%{\bf UNUSED REFERENCE:}  \ \ \
%R.~H.~Bing, {\it A decomposition of $E\sp 3$ into points and tame
%arcs such that the decomposition space is topologically different
%from $E\sp 3$}, Ann. of Math. (2) {\bf 65} (1957), 484--500.
%
%\bibitem{bing2}
%{\bf UNUSED REFERENCE:}  \ \ \
%R.~H.~Bing, {\it The cartesian product of a certain nonmanifold and
%a line is $E^4$}, Ann. of Math. (2) {\bf 70} (1959), 399--412.

\bibitem{Bing}
R. H. Bing,
{\sl Point-like decompositions of $E^3$},
Fund. Math. {\bf 50} (1961/1962), 431--453.

% \bibitem{Borsuk}
% K. Borsuk,
% {\it TITLE:} ,
% Mat. Sb. (N.S.) {\bf 1} (1936), 643--650

%\bibitem{Bredon}
%{\bf UNUSED REFERENCE:}  \ \ \
%G.~Bredon, {\it Generalized manifolds revisited}, ``Topology of
%Manifolds'' (J.D. Cantrell and C. H. Edwards, Eds.), Proceeding of
%the University of Georgia Topology of Manifolds Institute, Markham,
%Chicago, 1970, pp. 461--469.
%
%\bibitem{Bryant 1}
%{\bf UNUSED REFERENCE:}  \ \ \
%J.~L.~Bryant, {\it Euclidean space modulo a cell}, Fund. Math. {\bf
%63} (1968) 43--51.
%
%\bibitem{Bryant 2}
%{\bf UNUSED REFERENCE:}  \ \ \
%J.~L.~Bryant, {\it Euclidean $n$-space modulo an $(n-1)$-cell},
%Trans. Amer. Math. Soc. {\bf 179} (1973) 181--192.

%\bibitem{Bryant}
%J.~Bryant,
%{\it Homogeneous ENR's},
%Topology Appl. {\bf 27} (1987), 301--306.

%\bibitem{BFMW}
%J.~Bryant, S.~Ferry, W.~Mio and S.~Weinberger, {\it Topology of
%homology manifolds},
%Ann. of Math. (2) {\bf 143}:3 (1996), 435--467. % MR1394965 (97b:57017)

\bibitem{Busemann1}
H.~Busemann, {\it Metric methods in Finsler spaces and in the
foundations of geometry,} Ann. of Math. Studies  {\bf 8}, 
Princeton
University Press, Princeton, 1942.

\bibitem{Busemann2}
H.~Busemann, {\it On spaces in which two points determine a
geodesic,} Trans. Amer. Math. Soc. {\bf 54 } (1943), 171--184.

\bibitem{Busemann3}
H.~Busemann, {\it The Geometry of Geodesics}, Academic Press, New
York, 1955.

\bibitem{Cannon}
J.~W.~Cannon, {\it  The recognition problem: What is a topological
manifold?}, Bull. Amer. Math. Soc. {\bf 84} (1978), 832--866.

%\bibitem{Cannon 1}
%{\bf UNUSED REFERENCE:}  \ \ \
%J.~W.~Cannon, {\it The characterization of topological mani\-folds
%of dimension $\ge 5$}, Proc. Int. Congr. Math. Helsinki 1978,
%Academia Sci. Fennica, Helsinki, 1980, pp. 449--454.

%\bibitem{CaBrLa}
%J.~W.~Cannon, J.~L.~Bryant, and R.~C.~Lacher, {\it The structure of
%generalized manifolds having nonmanifold set of trivial dimension},
%Geometric Topology (Proc. Georgia Topology Conf., Athens, Ga.,
%1977), Academic Press, New York-London, 1979, pp. 261v-300.

%\bibitem{Cannon-Daverman}
%{\bf UNUSED REFERENCE:}  \ \ \
%J.~W.~Cannon and R.~J.~Daverman, {\it Cell-like decompositions
%arising from mismatched sewings: applications to 4-manifolds}, Fund.
%Math. {\bf 111} (1981), 211--233.

\bibitem{Cannon-Daverman2}
J.~W.~Cannon and R.~J.~Daverman, {\it  A totally wild flow}, Indiana
Univ. Math. J. {\bf 30}:3 (1981), 371--387.

\bibitem{CannonLO}
L. O. Cannon,
{\sl On a point-segment decomposition of $E^3$ defined by McAuley},
Proc. Amer. Math. Soc. {\bf 19} (1968), 624--630.

%\bibitem{Daverman 2}
%R.~J.~Daverman, {\it Singular regular neighborhoods and local
%flatness in codimension one,} Proc. Amer. Math. Soc. {\bf 57}
%(1976), 357--362.

\bibitem{Daverman S}
R.~J.~Daverman, {\it Products of cell-like decomposition}, Topology
Appl. {\bf 11} (1980), 121--139.

\bibitem{Daverman 1}
R.~J.~Daverman, {\it Detecting the disjoint disks property,} Pacific
J. Math. {\bf 93}:2 (1981), 277--298.

%\bibitem{Daverman}
%R.~J.~Daverman,
%{\it  Embedding phenomena based upon decomposition
%theory: locally spherical but wild codimension one spheres,}
%Proc. Amer. Math. Soc. {\bf 90}:1 (1984), 139--144.

\bibitem{Daverman book}
R.~J.~Daverman, {\it Decompositions of Manifolds}, Pure and Applied
Mathematics {\bf 124}, Academic Press, Inc., Orlando, FL, 1986.

\bibitem{Daverman-Halverson 2}
R.~J.~Daverman and D.~M.~Halverson, {\it Path concordances as
detectors of codimension-one manifold factors}, Exotic Homology
Manifolds, Oberwolfach 2003, Geom. Topol. Monogr. {\bf 9}, Geom.
Topol. Publ., Coventry, 2006, pp. 7--15.

%\bibitem{Daverman-Halverson}
%{\bf UNUSED REFERENCE:}  \ \ \
%R.~J.~Daverman and D.~M.~Halverson, {\it The cell-like approximation
%theorem in dimension $n = 5$,} Fund. Math. {\bf 197} (2007),
%81--121.

%\bibitem{DaRe 1}
%R.~J.~Daverman and D.~Repov\v s, {\it A new 3-dimensional shrinking
%theorem}, Trans.~Amer.~Math.~Soc. {\bf 315} (1989), 219--230.

%\bibitem{DaRe 2}
%R.~J.~Daverman and D.~Repov\v s, {\it  General position properties
%that characterize 3-manifolds}, Canad.~J.~Math. {\bf 44} (2)(1992),
%234--251.

%\bibitem{DaTh}
%R.~J.~Daverman and T.~L.~Thickstun,
%{\it The 3-manifold recognition problem},
%Trans. Amer. Math. Soc. {\bf 358}:12 (2006), 5257-v5270.
%
\bibitem{Daverman-Walsh}
R.~J.~Daverman and J.~J.~Walsh, {\it A ghastly generalized
$n$-manifold}, Illinois J. Math. {\bf 25}:4 (1981), 555--576.

%\bibitem{Dydak-Walsh}
%J.~Dydak and J.~J.~Walsh,
%{\it Sheaves that are
%locally constant with applications to homology manifold,}
%  ``Geometric Topology and shape Theory''
%(S. Marde\v si\' c and J. Segal, Eds.),
%Lecture Notes in Math. {\bf 1283},
%Springer-Verlag, Berlin and New York, 1987, pp. 65--87.

\bibitem{Eaton}
W. T. Eaton,
{\sl Applications of a mismatch theorem to decomposition spaces},
Fund. Math. {\bf 89}:3 (1975), 199--224.

%\bibitem{Edwards 1}
%{\bf UNUSED REFERENCE:}  \ \ \
%R.~D.~Edwards, {\it Demension theory, I}, Geometric Topology, Proc.
%Conf., Park City, Utah, 1974, Lecture Notes in Math. {\bf 438},
%Springer, Berlin, 1975, pp. 195--211.

\bibitem{Edwards 2}
R.~D.~Edwards, {\it The topology of manifolds and cell-like maps},
Proc. Int. Congr. of Math., Helsinki, 1978, Acad. Sci. Fennica,
Helsinki, 1980, pp. 111--127.

%\bibitem{Finsler}
%{\bf UNUSED REFERENCE:}  \ \ \
%P.~Finsler, {\it \"{U}ber Kurven und Fl\"{a}chen in allgemeinen
%R\"{a}umen}, Ph. D. Thesis, G\"{o}ttingen, 1918.

%\bibitem{Freedman}
%M.~H.~Freedman,
%{\it The topology of $4$-dimensional manifolds,}
%J. Diff. Geo. {\bf 17} (1982), 357--454.
%
%\bibitem{Freedman-Quinn}
%M.~H.~Freedman and F.~Quinn,
%{\it The Topology of $4$-Manifolds,}
%Princeton University Press, Princeton, 1990.

%\bibitem{Gillman-Martin}
%{\bf UNUSED REFERENCE:}  \ \ \
%D.~S.~Gillman and J.~M.~Martin, {\it Countable decompositions of
%$E^3$ into points and point-like arcs}, Notices Amer. Math. Soc.
%{\bf 10} (1963), 74--75.

\bibitem{Halverson 1}
D.~M.~Halverson, {\it Detecting codimension one manifold factors
with the disjoint homotopies property,} Topology Appl. {\bf 117}:3
(2002), 231--258.

\bibitem{Halverson 2}
D.~M.~Halverson, {\it $2$-ghastly spaces with the disjoint
homotopies property: the method of fractured maps,} Topology Appl.
{\bf 138}:1-3 (2004), 277--286.

\bibitem{Halverson 3}
D.~M.~Halverson, {\it Detecting codimension one manifold factors
with 0-stitched disks}, Topology Appl. {\bf 154}:9  (2007),
1993--1998.

\bibitem{HaRe}
D.~M.~Halverson and D.~Repov\v{s}, {\it The Bing-Borsuk and the
Busemann Conjectures}, Math. Comm. {\bf 13}:2 (2008), 163--184.

\bibitem{HaRe2}
D.~M.~Halverson and D.~Repov\v{s}, {\it  Detecting codimension one
manifold factors with topographical techniques}, Topology Appl. {\bf
156}:17 (2009), 2870--2880.

\bibitem{Bertinoro}
D.~M.~Halverson and D.~Repov\v{s},
{\em A survey on the Generalized R.~L.~Moore problem},
Proc. Conf. "Computational and Geometric Topology" (Bertinoro, Italy, June 17-19, 2010),
Atti Semin. Mat. Fis. Univ. Modena Reggio Emilia {\bf 58} (2011), 175--191.  

%\bibitem{Krakus}
%{\bf UNUSED REFERENCE:}  \ \ \
%B. Krakus, {\it Any $3$-dimensional $G$-space is a manifold,} Bull.
%Acad. Pol. Sci. {\bf 16} (1968), 737--740.

\bibitem{HW} W.~Hurewicz and H.~Wallman, {\it Dimension Theory},
Princeton Math. Series No. 4, Princeton Univ. Press, Princeton N.J.,
1941.

\bibitem{McAuley}
L. F. McAuley,
{\sl Another decomposition of $E^3$ into points and intervals},
Topology Seminar (Wisconsin, 1965),
Ann. of Math. Studies, No. 60, Princeton Univ.
Press, Princeton, N.J. 1966, pp. 33--51.

%\bibitem{Meyer}
%{\bf UNUSED REFERENCE:}  \ \ \
%D.~V.~Meyer, More decompositions of $E^3$ which are factors of
%$E^{n+1}$, Fund. Math. {\bf 67} (1970)49--65.

%\bibitem{Mitchell}
%W.~J.~R.~Mitchell,
%{\it Defining the boundary of a homology manifold},
%Proc. Amer. Math. Soc. {\bf 110} (1990), 509--513.

\bibitem{MiRe}
W.~J.~R.~Mitchell and D.~Repov\v s, {\it Topology of cell-like
mappings}, Proc. Conf. Diff. Geom. and Topol., Cala Gonone 1988,
Suppl. Rend. Fac. Sci. Nat. Univ. Cagliari, {\bf 58} (1988),
265--300.

\bibitem{Moore1}
R.~L.~Moore, {\it Concerning upper semicontinuous collections of
continua which do not separate a given set}, Proc. Nat. Acad. Sci.
{\bf 10} (1924), 356--360.

\bibitem{Moore2}
R.~L.~Moore, {\it Concerning upper semicontinuous collections of
continua,} Trans. Amer. Math. Soc. {\bf 27} (1925), 416--428.

%\bibitem{Quinn 1}
%F.~Quinn,
%{\it Ends of maps III:  dimensions $4$ and $5$},
%J. Diff. Geo. {\bf 17} (1982), 503--521.

%\bibitem{Quinn 2}
%F.~Quinn, {\it An obstruction to the resolution of homology
%manifolds}, Mich. Math. J. {\bf 34} (1987), 285--291.

%\bibitem{Raymond}
%F.~Raymond,
%{\it Separation and union theorems for
%generalized manifolds with boundary},
%Mich. Math. J. {\bf 7} (1960), 7--21.

\bibitem{Repovs1}
D.~Repov\v{s}, {\it The recognition problem for topological
manifolds}, Geometric and Algebraic Topology, J. Krasinkiewicz, S.
Spie\.{z}, and H. Toru\'{n}czyk, Eds., PWN, Warsaw 1986, pp.
77--108.

\bibitem{Repovs2}
D.~Repov\v{s}, {\it  Detection of higher dimensional topological
manifolds among topological spaces}, Giornate di Topologia e
Geometria Delle Variet\`{a}, Bologna 1990, M. Ferri, Ed., Univ.
degli Studi di Bologna 1992, pp. 113--143.

\bibitem{Repovs3}
D.~Repov\v{s}, {\it The recognition problem for topological
manifolds: A survey,} Kodai Math. J. {\bf 17}:3 (1994), 538--548.

%\bibitem{rourke-sanderson}
%{\bf UNUSED REFERENCE:}  \ \ \
%C.~P.~Rourke and B.~J.~Sanderson, {\it Introduction to
%Piecewise-Linear Topology}, Ergebnisse der Mathematik und ihrer
%Grenzgebiete {\bf 69}, Springer-Verlag, New York-Heidelberg, 1972.

%\bibitem{Thickstun 1}
%T.~L.~Thickstun,
%{\it  An extension of the loop theorem and resolution of
% generalized 3-manifolds with 0-dimensional singular set},
%Invent.~Math. {\bf 78} (1984), 161--222.
%
%\bibitem{Thickstun 2}
%T.~L.~Thickstun,
%{\it Resolutions of generalized 3-manifolds whose singular
%sets have general position dimension one},
%Topology~Appl. {\bf 138}:1-3 (2004),  61-v95.
%{\it  A further extension of the Loop theorem and resolutions of generalized 3-manifolds with 1-demensional singular set}, J. London Math. Soc.

%\bibitem{Thurston}
%{\bf UNUSED REFERENCE:}  \ \ \
%P. Thurston, {\it $4$-dimensional Busemann $G$-spaces are
%$4$-manifolds,} Diff. Geom. Appl. {\bf 6}:3 (1996), 245--270.

%\bibitem{Walsh}
%J.~J.~Walsh,
%{\it General position properties of generalized manifolds: a primer},
%  Proc.~Idaho State Topology Conf.,
%1985.

%\bibitem{Wilder}
%{\bf UNUSED REFERENCE:}  \ \ \
%R.~L.~Wilder, {\it Topology of Manifolds},
%Amer.~Math.~Soc.~Colloq.~Publ. {\bf 32}, Providence R.I., 1963.

\bibitem{Zemke} C.~F.~Zemke, {\it Dimension and decompositions},
Fund. Math. {\bf 95}:3 (1977), 157--165.

\end{thebibliography}

\documentclass{birkjour}
 \newtheorem{thm}{Theorem}[section]
 \newtheorem{cor}[thm]{Corollary}

 \theoremstyle{definition}
 \newtheorem{defn}[thm]{Definition}
 \theoremstyle{remark}
 \newtheorem{rem}[thm]{Remark}
 \newtheorem*{ex}{Example}

 \numberwithin{equation}{section}

\begin{document}

%
%
%
%
%
%
%
%
%

\title[Decompositions generating codimension one manifold factors]
 {Decompositions of $\mathbb{R}^n, n \geq 4,$ into convex sets 
  generate codimension 1 manifold factors}


\author[D. M. Halverson]{Denise M. Halverson}

\address{Department of Mathematics\\
Brigham Young University\\
Provo, UT 84602}

\email{halverson@math.byu.edu}



\author[D. Repov\v s]{Du\v san Repov\v s}
\address{Faculty of Education and\\
Faculty of Mathematics and Physics\\
University of Ljubljana\\
P.O. Box 2964\\
Ljubljana, Slovenia 1001}

\email{dusan.repovs@guest.arnes.si}


\subjclass{Primary 57N15, 57N75; Secondary 57P99, 53C70}

\keywords{Upper semicontinuous decomposition, convex set,
genera\-lized manifold,
cell-like resolution,
general position property,
codimension one mani\-fold factor,
Generalized Moore Problem}





\begin{abstract}
We show that if $G$ is an upper semicontinuous decomposition of
$\mathbb{R}^n$, $n \geq 4$, into convex sets, then the quotient
space $\mathbb{R}^n/G$ is a codimension one manifold factor.  In
particular, we show that $\mathbb{R}^n/G$ has the disjoint arc-disk
property.
\end{abstract}

\maketitle

\section{Introduction}

A space $X$ is said to be a \emph{codimension one manifold factor}
provided that $X \times \mathbb{R}$ is a manifold.  It is a long
standing unsolved problem as to whether all resolvable generalized
manifolds are codimension one manifold factors \cite{Daverman S}.
This is the so-called Product With a Line Problem and it is the
essence of the famous Generalized R.~L.~Moore Problem
\cite{Bertinoro, Moore1, Moore2}.

The Product With a Line Problem
speaks directly to one of the most fundamental questions in
geometric topology, which is how to recognize manifolds
\cite{Cannon, Edwards 2, Repovs1, Repovs2, Repovs3}. Because
manifolds have a rich structure which is useful to exploit in many
areas of mathematics and its applications, it is important to
recognize when one is dealing with a space that is a manifold.  One
notable example is the relevance of the Product With a Line Problem
to the famous Busemann Conjecture in metric geometry
\cite{Busemann1,Busemann2,Busemann3, HaRe}.

One might wonder even if a decomposition of $\mathbb{R}^n$ into
convex sets could  give rise to a decomposition space topologically
distinct from $\mathbb{R}^n$. This problem was investigated for
several years beginning with  Bing in the 1950's
\cite{Armentrout,Bing,CannonLO,Eaton, McAuley}. In 1970, Armentrout \cite{Armentrout} 
produced the first example of a decomposition of $\mathbb{R}^3$ into convex sets that yields a non-manifold.
Then in 1975, Eaton \cite{Eaton}
 demonstrated
 that a certain decomposition of
$\mathbb{R}^3$ into points and straight line segments, originally
proposed by Bing \cite{Bing}, is
indeed topologically distinct from
$\mathbb{R}^3$.  Hence, this type of complexity is
significant.  It should also be noted that there are no known 
examples of a non-manifold resulting from a decomposition of $\mathbb{R}^{n\geq4}$ into convex sets.

In this paper we shall investigate how the type of complexity
represented by decompositions of $\mathbb{R}^n$ into convex sets can
affect the classification of a decomposition space as a codimension
one manifold factor. We shall demonstrate that decompositions of
$\mathbb{R}^n$, $n \geq 4$, into convex sets are always codimension
one manifold factors. In particular, we shall show that such spaces
have a particularly strong general position property, the disjoint
arc-disk property.

\section{Preliminaries}

We briefly review some basic definitions and notations.  Recall that
a map $f: X \to Y$ is said to be \emph{proper} if whenever $C$ is a compact
subset of $Y$, then $f^{-1}(C)$ is compact.

There are various equivalent definitions of upper semicontinuous
decompositions \cite{Daverman book}, but the following will be the
most useful for our purposes:

\begin{defn}
A decomposition $G$ of $M$ into compact sets is said to be
\emph{upper semicontinuous}
(usc)
if and only if the associated
decomposition map $\pi: M \to M/G$ is a proper map.
\end{defn}

A compact subset $C$ of a space $X$ is said to be \emph{cell-like}
if for each neighborhood $U$ of $C$ in $X$, $C$ can be contracted to
a point
inside
$U$ \cite{MiRe}. A usc decomposition $G$ of $M$ is said to be
 \emph{cell-like} if each element $g \in G$ is
cell-like. A map $f: Y \to X$ is said to be
\emph{cell-like}
if
for each $x \in X$, $f^{-1}(x)$ is cell-like. A \emph{resolvable
generalized n-manifold} is an $n$-dimensional space $X$ that is the
image of a cell-like map $f:M \to X$ where $M$ is an $n$-manifold.

Convex sets are contractible, and hence they are cell-like. Thus, a
usc decomposition $G$ of $\mathbb{R}^n$ into convex sets is a
cell-like decomposition and the associated decomposition map $\pi:
\mathbb{R}^n \to \mathbb{R}^n/G$ is a cell-like map.  The fact that
$\mathbb{R}^n/G$ is finite-dimensional follows from a result of
Zemke (see \cite[Theorem 5.2]{Zemke}). Therefore, in this setting,
$\mathbb{R}^n/G$ is a resolvable generalized $n$-manifold.

For resolvable generalized manifolds, we have the following very
useful approximate lifting theorem, which follows from \cite[Theorem
17.1 and Corollary 16.12B]{Daverman book}:

\begin{thm} \label{ALT} Suppose that $G$ is a cell-like
decomposition of a manifold $M$,
with decomposition map $\pi: M \to
M/G $, and that
the quotient space
$M/G$ is finite-dimensional.
Then for any map $f: Z \to
M/G$  of a finite-dimensional compact polyhedron $Z$,
and any $\epsilon >0$,
there exists
a map $F:Z \to M$ such that $\pi F$
is an $\epsilon$-approximation of $f$.
\end{thm}


General position properties are very useful in detecting codimension
one manifold factors \cite{Daverman-Halverson 2,Halverson
1,Halverson 2,Halverson 3, HaRe2}.  For our results, we shall only
need to employ the following:

\begin{defn}
A space $X$ is said to have the \emph{disjoint arc-disk property}
(DADP) provided that any two maps $\alpha: I \to X$ and $f: D^2 \to
X$ can be approximated by maps with disjoint images, where $I$
denotes the unit interval and $D^2$ denotes a disk.
\end{defn}

\noindent The following theorem was demonstrated in \cite[Proposition
2.10]{Daverman 1}:

\begin{thm} \label{DADP thm} A resolvable generalized manifold having DADP is a codimension
one manifold factor.\end{thm}

Useful in discussions of the DADP is the local $0$-co-connectedness
property. A set $Z \subset X$ is said to have the \emph{local
$0$-co-connectedness property} ($0$-LCC) in $X$ if for every $z \in
Z \cap \text{Cl}(X-Z)$, each neighborhood $U$ of $z$ contains
another neighborhood $V$ of $z$ so that any two points in 
$V-Z$
are
path connected in 
$U-Z$.
Note that if $Z$ is nowhere dense in $X$,
then $Z =Z \cap \text{Cl}(X-Z)$.

The following theorem can be found in \cite[Corollary
26.2A]{Daverman book}:

\begin{thm} \label{0-LCC}
Each $k$-dimensional closed subset of a generalized $n$-manifold,
where $k \leq n-2$, is $0$-LCC.
\end{thm}

\noindent Since a $k$-dimensional closed subset of a generalized
$n$-manifold $X$, where $k \leq n-1$, is nowhere dense in $X$, we
have the following corollary: 

\begin{cor} \label{path}
If $Z$ is a $k$-dimensional closed subset of a generalized
$n$-manifold $X$, where $k \leq n-2$, then any path $\alpha: I \to
X$ can be approximated by a path  $\alpha': I \to X-Z$.
\end{cor}

\section{Main Results}

The main result of this paper is the following theorem:

\begin{thm}
Let $G$ be an upper semicontinuous decomposition of $\mathbb{R}^n$
into convex sets, where $n \geq 4$.  Then $\mathbb{R}^n/G$ is a
codimension one manifold factor.
\end{thm}

\noindent This theorem will follow immediately as a corollary of
Theorem \ref{DADP thm} and the following theorem:

\begin{thm} \label{DADP-4 thm}
Let $G$ be an upper semicontinuous decomposition of $\mathbb{R}^n$
into convex sets, where $n \geq 4$.  Then $\mathbb{R}^n/G$ has the
DADP.
\end{thm}

\begin{proof} Let $f: D^2 \to \mathbb{R}^n/G$ and $\varepsilon > 0$.
It follows from Corollary \ref{path} that it suffices to show that
there is an $\epsilon$-approximation $f': D^2 \to \mathbb{R}^n/G$ of
$f$ such that $f'(D^2)$ is $2$-dimensional.

Let $F: D^2 \to \mathbb{R}^n$ be a piecewise linear map, that is an
${\varepsilon}$-approximate lift of $f$. We shall show that $f' = \pi
F$ is then the
desired map.

Let $T$ denote a triangulation of $F(D^2)$. We claim that if
$\sigma$ is a $2$-simplex of $T$, then $f'(\sigma)$ is
$2$-dimensional. To see this, let $G_{\sigma}$ be the decomposition
induced over $\pi(\sigma)$, i.e. $G_{\sigma}$ is the decomposition
of $\mathbb{R}^n$ having as the only nontrivial elements, the
nontrivial elements of $G$ that meet $\sigma$.  Let $\omega:
\mathbb{R}^n \to \mathbb{R}^n/G_{\sigma}$ be the associated
decomposition map. Note that $\omega$ is necessarily a proper map,
being a decomposition induced over a closed set in the decomposition
space of a usc decomposition.

Let $P$ be the $2$-dimensional plane in  $ \mathbb{R}^n$ that contains $\sigma$. 
Let $\varpi$ denote the restriction of $\omega$ to $P$. Then $\varpi$ is also a
proper map.  Thus $\varpi$ determines a usc decomposition of the
plane into convex sets, elements that do not separate the plane. It
now follows from a classical result of Moore \cite{Moore1,Moore2},
that $\varpi$ is a near-homeomorphism onto its image. Thus
$\varpi(\sigma)$ is at most $2$-dimensional. 

But $\varpi(\sigma)$ is
homeomorphic to $\omega(\sigma)$, which in turn is homeomorphic to
$\pi(\sigma)$. Thus $\pi(\sigma)$ is at most $2$-di\-men\-sional
subset of $\mathbb{R}^n/G$. Hence $$f'(D^2) = \underset{\sigma \in
T^{(2)}}{\bigcup}\pi(\sigma)$$ is a $2$-dimensional subset of the
generalized $n$-manifold $\mathbb{R}^n/G$ \cite{HW}.
\end{proof}

\section{Conclusions}

As we have seen, the complexity represented by decompositions into
convex sets does not inhibit a decomposition space from being a
codimension one manifold factor. The fact that such spaces satisfy
the DADP is a pleasant result. 

It is well known that not all
codimension one manifold factors satisfy the DADP, and hence the
DADP is not a general position property that provides a
characterization of codimension one manifold factors. In fact, the
DADP condition is a
relatively weak tool for detecting codimension one
manifold factors, compared to other general position properties such
as:
\begin{itemize}
\item the disjoint homotopies property \cite{Halverson 1};
\item the plentiful $2$-manifolds property \cite{Halverson 1};
\item the $0$-stitched disks property \cite{Halverson 3};
\item the method of $\delta$-fractured maps \cite{Halverson 2}; and
\item the disjoint topographies (or disjoint concordance) property
\cite{Daverman-Halverson 2, HaRe2}. 
\end{itemize}
It is these stronger properties
that must be utilized to demonstrate that spaces such as the Totally
Wild Flow \cite{Cannon-Daverman2} and the Ghastly Spaces
\cite{Daverman-Walsh} are codimension one manifold factors.

In conclusion, we have demonstrated that we must look to other types
of complexities to realize a counterexample to the Generalized R.~L.~Moore
Problem, if  such an example does indeed exist.


\subsection*{Acknowledgment}
Supported by
the Slovenian Re\-search Agency grants
BI-US/11-12/023,
P1-0292-0101, J1-2057-0101 and
J1-4144-0101.
We thank the referee for several comments and suggestions.

\end{document}